17280(1-10%)^x = 11340
find x

Answer by Alan3354(69443)

17280(1-10%)^x = 11340
find x
-------------
17280(0.9)^x = 11340
0.9^x = 11340/17280 = 21/32
x*log(0.9) = log(21/32)
x = log(21/32)/log(0.9)
x =~ 3.99783
-------------
The (1 - 10%) is sloppy. Usually when using per cents, it's specificied to be 10% of something, not just 10%. This is a poorly written problem, which is why it was ignored for a long time.